Given an arithmetic sequence -11,2,15..., if T(m) is the first term that exceeds 200, find the value of m.
Answer by rothauserc(4718)

-11,2,15...
:
200 = -11 + 13(m-1)
:
13(m-1) = 211
:
13m -12 = 211
:
13m = 223
:
m = 17
:
for m = 17
:
T(17) = -11 +13(17-1) = 197
T(18) = -11 +13(18-1) = 210
:
*********************************************************
m = 18, T(18) = 210 > 200, is first term that exceeds 200
*********************************************************
: